MOA-II Obsevation Log Monthly Summary 2009/12
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | - | 0 | 0 | 0 |
LMC | 4560-4631 | 72 | 397 | 95200 |
SMC | 891-905 | 15 | 29 | 6450 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 87 | 426 | 101650 (28h00m850s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
20 | 19 | 18 | 13 | 12 | 19 | 68 | 65 | 63 | 22 | 19 | 19 |
Field13 | Field14 | Field15 | Field16 |
19 | 21 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
15 | 14 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2009/12/1: no observation / night time = 20040(sec)
2009/12/2: No observation / night time = 19860(sec)
2009/12/3: telescope analysis, change exposure time due to moon / night time = 19740(sec)
2009/12/4: image of chip2 looks like bias / night time = 19560(sec)
2009/12/5: chip2 is not working / night time = 19380(sec)
2009/12/6: chip2 is not working / night time = 19260(sec)
2009/12/7: No observation, Pumping on going / night time = 19140(sec)
2009/12/8: took FLAT with V-filter, chip2 is not working / night time = 18960(sec)
2009/12/9: No observation / night time = 18840(sec)
2009/12/10: chip2 is not working / night time = 18720(sec)
2009/12/11: No observation / night time = 18660(sec)
2009/12/12: No observation / night time = 18480(sec)
2009/12/13: No observation / night time = 18420(sec)
2009/12/14: No observation / night time = 18360(sec)
2009/12/15: chip2 is not working / night time = 18300(sec)
2009/12/16: observed LMC for the catalogue with offset / night time = 18240(sec)
2009/12/17: / night time = 18180(sec)
2009/12/18: no observation / night time = 18120(sec)
2009/12/19: No observation / night time = 18060(sec)
2009/12/20: / night time = 18060(sec)
2009/12/21: observed LMC and FLAT via V-band / night time = 18060(sec)
2009/12/22: CCD temperature is not enough cold. No observation / night time = 18060(sec)
2009/12/23: / night time = 18120(sec)
2009/12/24: No observation / night time = 18060(sec)
2009/12/25: No observation. / night time = 18120(sec)
2009/12/26: observed Dome flat / night time = 18180(sec)
2009/12/27: observed FLAT for gain. / night time = 18240(sec)
2009/12/28: No observation / night time = 18300(sec)
2009/12/29: No observation / night time = 18360(sec)
2009/12/30: / night time = 18480(sec)
2009/12/31: Exposure time is 150sec due to the full moon. / night time = 18600(sec)
TOTAL NIGHT TIME: 576960(sec)
OPERATING RATE: 20.6 %
GB: 0.0 %
LMC: 19.3 %
SMC: 1.3 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month